# Kinetix: Investigating the Training of General Agents through Open-Ended Physics-Based Control Tasks

### v1

@@ Abstract
While large models trained with self-supervised learning on offline datasets have shown remarkable capabilities in text and image domains, achieving the same generalisation for agents that act in sequential decision problems remains an open challenge. In this work, we take a step towards this goal by procedurally generating tens of millions of 2D physics-based tasks and using these to train a general reinforcement learning (RL) agent for physical control. To this end, we introduce Kinetix: an open-ended space of physics-based RL environments that can represent tasks ranging from robotic locomotion and grasping to video games and classic RL environments, all within a unified framework. Kinetix makes use of our novel hardware-accelerated physics engine Jax2D that allows us to cheaply simulate billions of environment steps during training. Our trained agent exhibits strong physical reasoning capabilities, being able to zero-shot solve unseen human-designed environments. Furthermore, fine-tuning this general agent on tasks of interest shows significantly stronger performance than training an RL agent tabula rasa. This includes solving some environments that standard RL training completely fails at. We believe this demonstrates the feasibility of large scale, mixed-quality pre-training for online RL and we hope that Kinetix will serve as a useful framework to investigate this further. [^1] We provide full code and models at https://kinetix-env.github.io.
@@ 2.3 RL in JAX
JAX (Bradbury et al., 2018) is a Python library for writing parallelisable code for hardware accelerators. While deep RL has traditionally been divided between environments on the CPU and models on the GPU (Mnih et al., 2015; Espeholt et al., 2018), JAX has facilitated the development of GPU-based environments (Lange, 2022; Rutherford et al., 2023; Nikulin et al., 2023; Matthews et al., 2024; Kazemkhani et al., 2024), allowing the entire RL pipeline to run on a hardware accelerator (Hessel et al., 2021). Through massive parallelisation and elimination of CPU-GPU transfer, this gives tremendous speed benefits (Lu et al., 2022). While UED has also followed this trend (Jiang et al., 2023; Coward et al., 2024), experiments have largely been confined to simple gridworlds, due to the lack of any suitable alternative (Garcin et al., 2024; Rutherford et al., 2024).
@@ 3.2 Kinetix: RL Environment Specification
Kinetix builds on Jax2D to create an environment for RL, which we now briefly outline.
@@
Reward To facilitate our goal of a general agent, we choose a simple yet highly expressive reward function that remains fixed across all environments. Each scene must contain a green shape and a blue shape—the goal is simply to make these two shapes collide, upon which the episode terminates with a reward of $+1$. Scenes can also contain red shapes, which, if they collide with the green shape, will terminate the episode with $-1$ reward. As demonstrated in Figure 1, these simple and interpretable rules allow for a large number of semantically diverse environments to be represented. To improve learning, we augment this sparse reward with an auxiliary dense reward signal, defined as $R^{d}_{t}=\kappa\left(d_{t}-d_{t+1}\right)$, where $d_{t}$ is the distance between the green and blue objects at timestep $t$ and $\kappa$ is a coefficient that we tune to ensure the dense signal does not dominate.
@@ 4 Experimental Setup
To select levels to train on, we use SFL (Rutherford et al., 2024), a state-of-the-art UED algorithm that regularly performs a large number of rollouts on randomly generated levels. It then selects a subset of these with high learnability and trains on them for a fixed duration before again selecting new levels. The main limitation of SFL, that it is only applicable to deterministic settings with a binary success reward, does not constrain us, as Kinetix satisfies both of these assumptions. We ran preliminary experiments using PLR (Jiang et al., 2021a; b) and ACCEL (Parker-Holder et al., 2022), but found that these approaches provided no improvements over DR (see Appendix L).
@@ 5 Zero-Shot Results
In Figure 3, we run SFL on the S, M and L environment sizes, respectively (see Appendix J for a per-level breakdown). In each case, the (random) training environments are of the corresponding size, and we use the corresponding holdout set (see Appendix E for a full listing) to evaluate the agent’s generalisation capabilities. We see that, in every case, the agent’s performance increases throughout training, indicating that it is learning a general policy that it can apply to unseen environments. For S, the agent very quickly learns a policy superior to the random policy, and is able to solve most of the hold out levels zero-shot. While the solve rate is lower on M, the agent can still zero-shot a number of unseen hand-designed environments. On the L environments, in which the agent is assessed on the most challenging holdout tasks, we see a very slow, and non-monotonic, performance increase. As well as being trained and tested on more complex levels, it seems that as the complexity increases, randomly generated levels are more likely to be unsolvable, reducing the proportion of useful data the agent can learn on. Overall, this result demonstrates that training an agent on a large set of mixed-quality levels can lead to general behaviour on unseen tasks.
@@ 6 Fine-Tuning Results
In this section we leave the zero-shot paradigm and investigate the performance of the general agent when given a limited number of samples to fine-tune on the holdout tasks. In particular, in Figure 5 we train a separate specialist agent for each level in the L holdout set, and compare this to fine-tuning a general agent (the same one used for Section 5.1). We plot the learning curves for four selected environments, as well as the aggregate performance over the entire holdout set. On three of these levels, fine-tuning the agent drastically outperforms training from scratch. In particular, for Mujoco-Hopper-Hard and Mujoco-Walker-Hard, the fine-tuned agent is able to competently complete these levels, whereas the tabula rasa agent cannot do so consistently. Notably, this is despite the fact that the pre-trained agent cannot solve these environments zero-shot. While the general trend is that fine-tuning beats training from scratch, we do see one case: Thruster-Large-Obstacles, where fine-tuning learns slower.
@@ 7 Related Work
Open-Ended Learning Kinetix also ties into the paradigm of open-ended learning (Soros & Stanley, 2014; Stanley, 2019; Sigaud et al., 2023; Hughes et al., 2024), in which a system continually generates new and novel artifacts. In the context of RL, this often means training inside a large, diverse distribution and having some method to adapt this distribution over time, with fields such as UED specifically focusing on the latter. While these methods hold the promise of generating novel and useful levels in an open-ended manner, the environments used in their experiments are often very constrained in what they can represent (Wang et al., 2019; Dennis et al., 2020; Wang et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022). As we have shown, in a significantly more diverse task space, these methods tend to fail.
A recent work with a similar vision to Kinetix is Autoverse (Earle & Togelius, 2024), where an agent acts inside a cellular automata (CA) based gridworld. By changing the underlying rules of the CA, a large diversity of different levels can be represented in this system. Relatedly, Sun et al. (2024) use prior knowledge in the form of Large Language Models to generate code for video games and robotic simulations, to train RL agents on. Powderworld (Frans & Isola, 2023) instead creates a expressive environment based on different types of elements interacting in a sandbox environment. Other notable work that aims to use open-ended discovery to train generalist agents include Voyager (Wang et al., 2024), Jarvis-1 (Wang et al., 2023) and Optimus-1 (Li et al., 2024b). These are more focused on long-horizon planning, the self discovery of new tasks to perform, and use Minecraft as their domain with prior knowledge in the form of a large language model.
Perhaps the works most similar to ours is the highly impactful XLand line of research (Team et al., 2021; 2023). XLand defines a large and diverse distribution of levels inside a 3D physics simulation, with an embodied agent (or set of agents) required to fulfill some specified goal in the environment. Similar to us, XLand agents train on a procedurally generated set of levels and are then assessed on human-designed holdout levels. We see the main differences to Kinetix being the expressivity of the tasks and the public state of the work. In particular, we would subjectively claim that Kinetix, through representation of almost any conceivable 2D rigid-body physics problem, has a more expressive universe of tasks. While XLand also employs a physics engine, all the tasks are constrained to homogeneous agents acting in the world, potentially limiting the scope of tasks—it is not clear, for instance, how one would represent any of the holdout environments in Figure 1 in XLand. Lastly, we note that the XLand environment has not been made open-source, limiting its use for future research. XLand-Minigrid (Nikulin et al., 2023) provides a fast, open-source version of XLand, however in doing so it simplifies the environment into a gridworld.
@@ 8 Discussion and Future Work
We believe Kinetix is a uniquely diverse, fast and open-ended environment, placing it well as a foundation to study open-ended RL, including large-scale online pre-training for general RL agents. In stark contrast to many other benchmarks used for open-ended learning (Wang et al., 2019; 2020; Parker-Holder et al., 2022; Chevalier-Boisvert et al., 2023; Rutherford et al., 2024), Kinetix represents a large space of semantically diverse tasks, instead of just variations on a single task. This presents a challenge for future environment design research that can intelligently generate levels (Dennis et al., 2020), rather than just filtering from a predefined distribution. We also believe Kinetix is an excellent framework for investigating issues in agent training like network capacity (Obando-Ceron et al., 2024), plasticity loss (Igl et al., 2020; Berariu et al., 2021; Dohare et al., 2021; Sokar et al., 2023), lifelong learning (Kirkpatrick et al., 2017) and multi-task learning (Sodhani et al., 2021; Hafner, 2021; Benjamins et al., 2023).

### v2

@@ Abstract
While large models trained with self-supervised learning on offline datasets have shown remarkable capabilities in text and image domains, achieving the same generalisation for agents that act in sequential decision problems remains an open challenge. In this work, we take a step towards this goal by procedurally generating tens of millions of 2D physics-based tasks and using these to train a general reinforcement learning (RL) agent for physical control. To this end, we introduce Kinetix: an open-ended space of physics-based RL environments that can represent tasks ranging from robotic locomotion and grasping to video games and classic RL environments, all within a unified framework. Kinetix makes use of our novel hardware-accelerated physics engine Jax2D that allows us to cheaply simulate billions of environment steps during training. Our trained agent exhibits strong physical reasoning capabilities in 2D space, being able to zero-shot solve unseen human-designed environments. Furthermore, fine-tuning this general agent on tasks of interest shows significantly stronger performance than training an RL agent tabula rasa. This includes solving some environments that standard RL training completely fails at. We believe this demonstrates the feasibility of large scale, mixed-quality pre-training for online RL and we hope that Kinetix will serve as a useful framework to investigate this further. [^1] We provide full code and models at https://kinetix-env.github.io.
@@ 2.3 RL in JAX
JAX (Bradbury et al., 2018) is a Python library for writing parallelisable code for hardware accelerators. While deep RL has traditionally been divided between environments on the CPU and models on the GPU (Mnih et al., 2015; Espeholt et al., 2018), JAX has facilitated the development of GPU-based environments (Lange, 2022; Koyamada et al., 2023; Rutherford et al., 2023; Nikulin et al., 2023; Matthews et al., 2024; Kazemkhani et al., 2024; Bonnet et al., 2024; Pignatelli et al., 2024), allowing the entire RL pipeline to run on a hardware accelerator (Hessel et al., 2021). Through massive parallelisation and elimination of CPU-GPU transfer, this gives tremendous speed benefits (Lu et al., 2022). While UED has also followed this trend (Jiang et al., 2023; Coward et al., 2024), experiments have largely been confined to simple gridworlds, due to the lack of any suitable alternative (Garcin et al., 2024; Rutherford et al., 2024).
@@ 3.2 Kinetix: RL Environment Specification
Kinetix builds on Jax2D to create an environment for RL, which we now briefly outline. See Appendix C for further information.
@@
Reward To facilitate our goal of a general agent, we choose a simple yet highly expressive reward function that remains fixed across all environments. Each scene must contain a green shape and a blue shape—the goal is simply to make these two shapes collide, upon which the episode terminates with a reward of $+1$. Scenes can also contain red shapes, which, if they collide with the green shape, will terminate the episode with $-1$ reward. As demonstrated in Figure 1, these simple and interpretable rules allow for a large number of semantically diverse environments to be represented. To improve learning, we augment this sparse reward with an auxiliary dense reward signal, defined as $R^{d}_{t}=\kappa\left(d_{t}-d_{t+1}\right)$, where $d_{t}$ is the distance between the green and blue objects at timestep $t$ and $\kappa$ is a coefficient that we tune to ensure the dense signal does not dominate. We note that Kinetix could be run with many other reward formulations (Andrychowicz et al., 2017; Frans et al., 2024), which we leave to future work.
@@ 4 Experimental Setup
To select levels to train on, we use SFL (Rutherford et al., 2024), a state-of-the-art UED algorithm that regularly performs a large number of rollouts on randomly generated levels. It then selects a subset of these with high learnability and trains on them for a fixed duration before again selecting new levels. SFL filters out all unsolvable levels, as the success rate (and therefore also learnability) is zero. The main limitation of SFL, that it is only applicable to settings with deterministic transition dynamics and binary rewards, does not constrain us, as Kinetix satisfies both of these assumptions. We ran preliminary experiments using PLR (Jiang et al., 2021a; b) and ACCEL (Parker-Holder et al., 2022), but found that these approaches provided no improvements over DR (see Appendix L).
@@ 5 Zero-Shot Results
In Figure 3, we run SFL on the S, M and L environment sizes, respectively (see Appendix J for a per-level breakdown). In each case, we train on randomly-generated environments of the corresponding size, and we use the corresponding holdout set (see Appendix E for a full listing) to evaluate the agent’s generalisation capabilities. We see that, in every case, the agent’s performance increases throughout training, indicating that it is learning a general policy that it can apply to unseen environments. For S, the agent very quickly learns a policy superior to the random policy, and is able to solve most of the hold out levels zero-shot. While the solve rate is lower on M, the agent can still zero-shot a number of unseen hand-designed environments. On the L environments, in which the agent is assessed on the most challenging holdout tasks, we see a very slow, and non-monotonic, performance increase. As well as being trained and tested on more complex levels, it seems that as the complexity increases, randomly generated levels are more likely to be unsolvable, reducing the proportion of useful data the agent can learn on. Overall, this result demonstrates that training an agent on a large set of mixed-quality levels can lead to general behaviour on unseen tasks. See Appendix K for more detailed results.
@@ 6 Fine-Tuning Results
In this section we leave the zero-shot paradigm and investigate the performance of the general agent when given a limited number of samples to fine-tune on the holdout tasks. In particular, in Figure 5 we train a separate specialist agent for each level in the L holdout set, and compare this to fine-tuning a general agent (the same one used for Section 5.1, trained for 5B timesteps on random M levels.). We plot the learning curves for four selected environments, as well as the aggregate performance over the entire holdout set. On three of these levels, fine-tuning the agent drastically outperforms training from scratch. In particular, for Mujoco-Hopper-Hard and Mujoco-Walker-Hard, the fine-tuned agent is able to competently complete these levels, whereas the tabula rasa agent cannot do so consistently. Notably, this is despite the fact that the pre-trained agent cannot solve these environments zero-shot. While the general trend is that fine-tuning beats training from scratch, we do see one case: Thruster-Large-Obstacles, where fine-tuning learns slower.
@@ 7 Related Work
Open-Ended Learning Kinetix also ties into the paradigm of open-ended learning (Soros & Stanley, 2014; Stanley, 2019; Sigaud et al., 2023; Hughes et al., 2024), in which a system continually generates new and novel artifacts. In the context of RL, this often means training within a large and diverse distribution and applying some method (e.g., UED) to adapt this distribution over time. While these methods hold the promise of generating novel and useful levels in an open-ended manner, the environments used in their experiments are often very constrained in what they can represent (Wang et al., 2019; Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022). As we have shown, in a significantly more diverse task space, these approaches tend to fail.
A recent work with a similar vision to Kinetix is Autoverse (Earle & Togelius, 2024), where an agent acts inside a cellular automata based gridworld, where changing the underlying rules can lead to many diverse levels. Relatedly, Sun et al. (2024) use prior knowledge in the form of large language models to generate simulation code to train RL agents in. Powderworld (Frans & Isola, 2023) instead creates an expressive environment based on different types of elements interacting in a sandbox environment. Other notable work that aims to use open-ended discovery to train generalist agents include Voyager (Wang et al., 2024), Jarvis-1 (Wang et al., 2023) and Optimus-1 (Li et al., 2024b). These are more focused on long-horizon planning, the self discovery of new tasks to perform, and use Minecraft as their domain with prior knowledge in the form of a large language model.
Perhaps the work most similar to ours is the highly impactful XLand line of research (Team et al., 2021; 2023). XLand defines a large and diverse distribution of levels inside a 3D physics simulation, with an embodied agent (or set of agents) required to fulfil some specified goal Similar to us, agents train on procedurally generated levels and are assessed on human-designed holdout levels. We see the main differences to Kinetix being the expressivity of the tasks and the public state of the work. In particular, we subjectively claim that Kinetix, through the representation of almost any conceivable 2D rigid-body physics problem, has a more expressive universe of tasks. While XLand also employs a physics engine, all the tasks are constrained to homogeneous agents acting in the world, potentially limiting its scope—it is not clear, for instance, how one would represent any of the holdout environments in Figure 1 in XLand. Lastly, we note that XLand’s source code is unavailable, limiting its use for future research. Although XLand-Minigrid (Nikulin et al., 2023) provides a fast, open-source version of XLand, it simplifies the environment into a gridworld.
@@ 8 Discussion and Future Work
We believe Kinetix is a uniquely diverse, fast and open-ended environment, placing it well as a foundation to study open-ended RL, including large-scale online pre-training for general RL agents. In stark contrast to many other benchmarks used for open-ended learning (Wang et al., 2019; Chevalier-Boisvert et al., 2023; Rutherford et al., 2024), Kinetix represents a large space of semantically diverse tasks, instead of just variations on a single task. This presents a challenge for future environment design research that can intelligently generate levels (Dennis et al., 2020), rather than just filtering from a predefined distribution. We also believe Kinetix is an excellent framework for investigating issues in agent training such as network capacity (Obando-Ceron et al., 2024), plasticity loss (Igl et al., 2020; Berariu et al., 2021; Sokar et al., 2023), lifelong learning (Kirkpatrick et al., 2017) and multi-task learning (Sodhani et al., 2021; Hafner, 2021; Benjamins et al., 2023).
@@ 9 Conclusion
Acknowledgements
We would like to thank Thomas Foster, Alex Goldie, Matthew Jackson, Sebastian Towers, Andrei Lupu and our anonymous reviewers for insightful discussions and valuable feedback that aided the development of this project and the production of the manuscript. This work was supported by UK Research and Innovation and the European Research Council, selected by the ERC, and funded by the UKRI [grant number EP/Y028481/1]. We also thank the authors of the game Incredibots, which served as an initial inspiration for the development of Kinetix.
@@ Acknowledgements
  • Appendix N compares the learnability of chosen vs randomly sampled environments over the course of training.
  • Appendix O Ablates the observation and action spaces of Kinetix.
  • Finally, Appendix Q briefly investigates lifelong learning aspects of the general agent.
@@ Appendix B Jax2D Speed Results
For actual runtimes, training the generalist agent for 1 billion timesteps on a single L40S took around 7 hours for S, 9 hours for M and 14 hours for L. Training on such a large number of timesteps is indeed nontrivial, but JAX and our Jax2D engine makes it feasible. This could further be sped up by using multiple GPUs in parallel.
@@ Appendix I Specialist Results
Appendix N Learnability Over Training
In Figure 29, we plot the learnability of the training levels, and the larger set of random levels these are sampled from. Overall, we find that random levels tend to have a low learnability, whereas the top 1024 levels consistently has high learnability throughout training.
Figure 29: The learnability (scaled to between 0 and 1) over training for each of the environment sizes. We compute learnability for 12288 randomly-generated levels (shown in orange), and select the top 1024 of these (the learnability of this top subset is shown in blue).
Appendix O Alternate Observation and Action Spaces
In Figure 30, we consider a multi-task setting, where we train agents on the holdout tasks with each combination of MultiDiscrete/Continuous action space and Pixels/Symbolic-Entity observation space. We find that MultiDiscrete actions outperform Continuous actions, and that Entity slightly outperforms Pixels (in addition to being significantly faster to run). This experiment validates our decision to use Entity observations and MultiDiscrete actions for our main experiments.
Figure 30: We compare the different observation and action spaces on a multi-task setting. Here we trained agents on all 74 holdout tasks, and show aggregate performance on this set of tasks. We plot mean and standard deviation over five seeds.
Appendix P Cross-Embodiment Learning
In Figure 31, we consider cross-embodiment learning, where we train a single agent to control all 7 of the Mujoco recreations. We compare this against agents trained individually for each task. For fairness, we allocate more samples to the single agent (500M vs 100M). We find that the single agent is able to competently control all morphologies, although it is less sample efficient when considering only a single task. On some tasks (e.g. MuJoCo-Walker) we see improved learning from co-training with other morphologies.
Figure 31: Comparing the performance of agents individually trained against one jointly trained on the recreations of Mujoco tasks. Combined indicates the agent trained jointly, and all plots show mean and shade standard deviation over 5 seeds. We note that the x-axis measures the total number of timesteps, i.e., for the Combined line, this includes all morphologies.
Appendix Q Lifelong Learning
In Figure 32, we plot a single training run where we first train an agent on random levels from the S distribution for 5B timesteps. We then change this and train the agent on random M levels for 1B timesteps and finally train it again on random S levels for 1B timesteps. We plot the performance of the agent on the heldout set of levels for both the S and M size separately. As expected, training on S initially slightly improves performance on the M set of holdout levels. Then, training for 1B timesteps on M improves performance by a larger margin. Going back to training on random S levels reduces the performance on the M holdout set. This indicates a level of forgetting or plasticity loss in the agent.
Figure 32: A single run’s training, where we first train the agent on S for 5B timesteps, then transition to M for 1B and finally train on S again for 1B. We plot the aggregate evaluation performance on the S set of holdout levels on the left and the M holdout levels on the right.